OpenStudy (anonymous):

I need help with this problem.. -2(v - 7) - 6v = 8(4 - v) - 3

OpenStudy (anonymous):

-2v + 14 - 6v = 32- 8v - 3 then add like terms.

OpenStudy (anonymous):

After combining like terms, would I get 2v = 17 ? Do I have to divide by 2 on both sides? To get v alone?

OpenStudy (anonymous):

v=15 -2v-6v+8v will give you 0v and 32-3-14 will give you 15

OpenStudy (anonymous):

Oh I see, thank you very much!